Systems of Inequalities:
A dry cleaner cleans neckties for $3.25 each; however, silk neckties are cleaned for $3.60 each. A customer brings in 17 neckties of various materials. The total bill is $58.75. How many silk ties did the customer bring in?

Answer by checkley77(12569)

3.60x+3.25(17-x)=58.75
3.60x+55.25-3.25x=58.75
.35x=58.75-55.25
.35x=3.50
x=3.50/.35
x=10 silk ties were cleaned.
17-10=7 non silk ties.
Proof:
3.60*10+3.25*7=58.75
36.00+22.75=58.75
58.75=58.75